OpenStudy (cherie_magee):

Find the 31st term of the following sequence. 9, 15, 21, ... 186 189 195

OpenStudy (goldphenoix):

What's the increase from 9 to 15?

OpenStudy (goldphenoix):

Are you there, @cherie_magee

OpenStudy (cherie_magee):

yeah...an 6 @GoldPhenoix

OpenStudy (goldphenoix):

Good. What was the increase from 15 to 21?

OpenStudy (cherie_magee):

it's all 6 i just think thats toooo much to do

OpenStudy (cherie_magee):

@GoldPhenoix

OpenStudy (texaschic101):

an = a1 + (n-1)d n (the number of the term you want to find) = 31 a1 (the first term) = 9 d (the common difference) = 6 now sub that into your formula..... a31 = 9 + (31 - 1) * 6 a31 = 9 + (30 * 6) I believe you can finish it from here :)

OpenStudy (cherie_magee):

LOL THANK YA SUGAR FOOT ;* @texaschic101

OpenStudy (texaschic101):

no problem :)